### v5

# Disentangling the Complex Multiplexed DIA Spectra in De Novo Peptide Sequencing

Abstract
Data-Independent Acquisition (DIA) was introduced to improve sensitivity to cover all peptides in a range rather than only sampling high-intensity peaks as in Data-Dependent Acquisition (DDA) mass spectrometry. However, it is not very clear how useful DIA data is for de novo peptide sequencing as the DIA data are marred with coeluted peptides, high noises, and varying data quality. We present a new deep learning method DIANovo, and address each of these difficulties, and improves the previous established system DeepNovo-DIA by from 34% to 108%, averaging 50%, for amino acid recall, and by from 32% to 83%, averaging 57%, for peptide recall, by equipping the model with a deeper understanding of coeluted DIA spectra. This paper also provides criteria about when DIA data could be used for de novo peptide sequencing and when not to by providing a comparison between DDA and DIA, in both de novo and database search mode. We find that while DIA excels with narrow isolation windows on older-generation instruments, it loses its advantage with wider windows. However, with Orbitrap Astral, DIA consistently outperforms DDA due to narrow window mode enabled. We also provide a theoretical explanation of this phenomenon, emphasizing the critical role of the signal-to-noise profile in the successful application of de novo sequencing.
1 Introduction
De novo peptide sequencing plays a crucial role in proteomics, enabling the identification of novel peptides, post-translational modifications, and mutations absent from existing protein databases[1][2][3][4]. This capability is crucial for personalized immunotherapy, guiding targeted treatments by identifying unique neoantigens. It is also essential for studying species with unsequenced genomes, where database searches are not feasible [5] [6] [7].
Compared to database-driven methods, de novo sequencing offers the advantage of discovering new peptide sequences independently of pre-existing data, enables capturing the full diversity of proteomes, especially in complex and dynamic biological systems [7] [8] [9]. In traditional Data-Dependent Acquisition (DDA) methods, where abundant peptides are selected to fragment[10] [11], deep learning-based approaches such as DeepNovo[2], Casanovo[12], and GraphNovo[13] have significantly improved performance, making them valuable and efficient tools for biological research.
Despite their promising results, traditional DDA methods suffer from limitations such as biased sampling and inconsistent detection of low-abundance peptides, leading to incomplete proteome coverage [11] [14] [15]. Data-Independent Acquisition (DIA) [15][16] addresses these limitations by fragmenting all ionized peptides within a predefined mass range, providing a more comprehensive and unbiased snapshot of the proteome [14].
DIA data fundamentally differ from DDA data in that DIA produces highly multiplexed spectra containing fragment ions from multiple coeluting peptides, where fragments of multiple peptides coexist in the same spectrum [17] [18]. However, these characteristics come with significant challenges for de novo peptide sequencing. DIA spectra are inherently noisier due to the simultaneous fragmentation of all peptides, complicating the distinction between signal peaks and noise. The concurrent fragmentation leads to overlapping fragmentation patterns, making it difficult to assign fragment ions to specific peptides. The mixed signals from multiple coeluting peptides increase the complexity of spectral interpretation compared to DDA, where each spectrum typically associates with a single precursor ion. Furthermore, the need to process and analyze these complex, noisier spectra increases computational demands. In addition, DIA provides chromatogram information that offers a temporal profile of peptide ions, and the similarities between these chromatograms could help trace the source of the corresponding ion.
Recent advancements such as DeepNovo-DIA [19], Transformer-DIA [20], and Cascadia [21] have harnessed the power of deep learning to address the challenges inherent to de novo sequencing in DIA, while approaches such as PepNet [22] have sought to develop unified models capable of handling both DDA and DIA data modalities.. These models employ neural networks to analyze precursor and fragment ions across multiple dimensions, including mass-to-charge ratio (m/z), retention time, and intensity, thus effectively handling the complexity of highly multiplexed spectra. For instance, DeepNovo-DIA integrates Ion-CNN and Spectrum-CNN  [23] with a long short-term memory (LSTM)  [24] network to capture the three-dimensional structure of fragment ions and their interrelationships. Similarly, Transformer-DIA adopts a Transformer-based  [25] architecture, using an encoder-decoder framework to predict peptide sequences by iteratively generating each subsequent amino acid based on prior outputs. Despite their efficacy, both models are constrained by their focus on encoding only a portion of the spectrum, potentially missing important signals related to future amino acids.
We propose DIANovo, a framework specifically designed to address the complexities of DIA data introduced by coelution. Concretely, we first set to tackle the sheer size of highly-multiplexed DIA spectrum. Our encoder encodes the spectrum graph, like GraphNovo[13]. To reduce memory consumption, we implements an automated edge generation process, which primarily focuses on learning the relative mass differences between graph nodes while the inter-node amino acid information is learned in a task-specific manner. To further reduce computational overhead, we leverage FlashAttention 2 [26], a linear-memory Transformer architecture designed for efficient computation. We also integrate dilated convolutional neural networks (CNNs)  [27] [28] to directly process chromatograms, enabling efficient encoding of time series information. Time-series embeddings are fed into the Transformer encoder, where the self-attention mechanism explicitly learns the similarities of the chromatograms, helping us to distinguish signal ions from those of coeluting peptides and noise. Moreover, we introduce a coelution-aware pretraining step that further improves model performance by incorporating coeluting peptides alongside the target peptide. This involves pretraining a model to predict ion types from coeluting peptides, with the resulting embeddings used as features in subsequent training stages. The coelution information helps the model better differentiate between signal and noise, leading to more accurate target peptide predictions. To the best of our knowledge, this is the first work to leverage coeluting peptide information in de novo sequencing, addressing a gap overlooked by previous methods.
Additionally, our study seeks to assess the realistic performance of DIA de novo sequencing. We report amino acid and peptide recalls compared with DIA-NN [29] search outcomes, considering only peptides unique to the test set. Our findings indicate that de novo peptide sequencing using DIA on older-generation instruments like the Q Exactive[30] or Fusion [31] is suboptimal, yielding relatively lower peptide recall. However, the Orbitrap Astral [32] significantly outpaces older-generation in acquisition speed, enabling the use of narrow-window data-independent acquisition (nDIA). The Astral narrow-window DIA method [32], which offers more consistent fragmentation patterns and fewer missing fragmentations, even in case of high coelution level, yields better results.
To better understand these challenges, we developed a theoretical framework that explains how the balance between signal enhancement and noise accumulation in different acquisition methods affects de novo peptide sequencing performance. Specifically, we analyze the signal and noise characteristics of DDA, older-generation DIA, and Astral DIA, and their impact on the efficacy of peptide matching algorithms like XCorr.[33] [34] Our model reveals that older-generation DIA introduces substantial noise that outweighs the marginal gain in signal peaks—diminishing de novo sequencing performance—while Astral DIA provides a disproportionate increase in signal peaks despite higher noise levels. This increase effectively enhances peptide identification by improving the confidence of peptide-spectrum matches. This theoretical insight underscores the critical importance of optimizing signal-to-noise ratios in mass spectrometry data to improve de novo sequencing outcomes.
Our extensive experiments across various datasets, including older-generation and Astral data, highlight the robustness and effectiveness of our proposed method. We demonstrated that our model consistently outperformed the baselines DeepNovo-DIA and Transformer-DIA. These findings underscore the potential of our method for robust and accurate de novo peptide sequencing in the challenging DIA setting.
Finally, we address a key question: can de novo sequencing in DIA mode detect more peptides than DDA mode? We present a comparison between DDA and DIA data acquired from the same biological sample, demonstrating that with older-generation mass spectrometers, DIA surpasses DDA in peptide detection when using smaller isolation windows. However, as the isolation windows widen, DIA loses this advantage and falls behind DDA. With Astral data, DIA can consistently detect more peptides than DDA because of the narrow DIA method enabled, showcasing the superior performance of next-generation mass spectrometry.
2 Results
Our experiments comprehensively evaluated the performance and robustness of our de novo peptide sequencing algorithm across various datasets and conditions. We found that older-generation mass spectrometers exhibit relatively lower peptide recall, while Astral data showed better results due to improved fragment ion coverage. Our model consistently outperformed the baselines DeepNovo-DIA and Transformer-DIA model across multiple datasets, highlighting its superior capability in complex DIA settings. We also include a comparison with a recent state-of-the-art model, Cascadia, demonstrating the superior performance of our approach.
Furthermore, in the comparison of DDA and DIA on the Hela [35] (older-generation) and PXD046453 [32] Hek293T (Astral) dataset, DIA demonstrated advantage in peptide detection when isolation window is small, with our de novo sequencing algorithm identifying additional peptides that DDA missed. These findings collectively underscore the robustness, versatility, and superior performance of DIA in diverse and challenging proteomics scenarios. These findings provide valuable insights into the proper scenarios for applying DIA de novo sequencing and highlight DIA’s capability to extend identification beyond DDA. Finally, we provide an ablation study justifying our model design choices.
2.1 De Novo Performance on Older-Generation Data
We compared our model’s performance with the baselines across several datasets [36] [37] [35] in Figure 1, where peptide recall is defined as the proportion of peptides whose entire sequences exactly match the database search results, without any mismatches or errors. On most datasets, our model demonstrated significantly better performance than the baselines, however the results are relatively lower across the board compared to DDA levels [2] [13], indicating that there is a significant gap between database and de novo identification performance in older-generation mass spectrometers. Our amino acid recall is on average 60% higher than DeepNovo-DIA, while peptide recall being 53% higher. Comparing to Transformer DIA, our amino acid recall and peptide recall is 32% and 24% higher respectively.
((a)) Amino acid recall
((b)) Peptide recall
((c)) Amino acid precision
Figure 1: Amino acid recall (a), peptide recall (b), and amino acid precision (c) of our method vs DeepNovo-DIA and Transformer-DIA, training sequences excluded from test set, on various older-generation datasets, measured over 10,000 randomly selected peptide precursors per dataset.
Although our method shows greater ability beyond the baselines, the results indicate that further refinements are needed to achieve satisfactory accuracy in de novo peptide sequencing in older-generation instruments.
2.2 Performance on Orbitrap Astral Data
Our experiments on Astral data demonstrate superior de novo performance compared to Orbitrap data. To ensure robustness and eliminate bias, we excluded training sequences from the test sets. Specifically, for human datasets (PXD046453 Hek293T, PXD046283 Hek293T, PXD046471 Clinical Sample), the model was trained on Yeast KO data, and for the Yeast KO dataset, the model was trained on the PXD046453 Hek293T dataset.
Our findings indicate that Astral data outperforms older-generation data in de novo peptide sequencing. This superior performance can be attributed to better fragment ion coverage in the Astral data, resulting in fewer missing fragments during the sequencing process. The performance differences are visually represented in the Figure 2, comparing our model’s performance against the baseline across the different datasets.
((a)) Amino acid recall
((b)) Peptide recall
((c)) Amino acid precision
Figure 2: Amino acid recall (a), peptide recall (b), and amino acid precision (c) of our method vs baselines on various Astral datasets.
The performance of our model on the Astral datasets shows a marked improvement over the baselines. On average we can expect a 43%/60% increase of amino acid / peptide recall with our methods compared to DeepNovo-DIA, or a 27%/47% improvement compared to Transformer-DIA. Furthermore, Astral data delivers a significant boost to de novo performance, affirming the effectiveness of DIA de novo sequencing in such system.
2.3 Comparison of Peptide Detection by DDA and DIA
((a)) 5 Th DIA window
((b)) 10 Th DIA window
((c)) 20 Th DIA window
Figure 3: Venn diagram, comparison of peptide identification under DDA or DIA mode, with Orbitrap Q Exactive (older-generation), where blue and orange circles refer to number of peptides identified in database search, while pink and green circles refer to number of peptides identified in de novo mode, under DDA and DIA respectively.
((a)) 2.5ms DIA Data
((b)) 3.5ms DIA Data
Figure 4: Venn diagram, comparison of peptide identification Under DDA or DIA mode, with Orbitrap Astral.
In this experiment, we conducted a detailed comparison of the number of peptides detected by de novo sequencing using both Data-Dependent Acquisition (DDA) and Data-Independent Acquisition (DIA) modes, utilizing both older-generation mass spectrometers and the newer Orbitrap Astral model. The goal was to understand how these acquisition strategies perform across different isolation windows, particularly in the context of de novo peptide identification.
For older-generation instruments, we adopted the Hela dataset [35], analyzing peptides identified from the same biological sample in the DDA and DIA modes with varying DIA isolation windows (5 Th, 10 Th, and 20 Th). In the DDA experiments, we performed a database search using the PEAKS[38] DB search engine and employed PointNovo [39] for de novo peptide sequencing. For DIA, we utilized the DIA-NN[29] search engine for database search and applied our de novo sequencing methods.
The results (shown in Figure 3) reveal a clear relationship between the isolation window size and the efficacy of peptide detection. When the isolation window is narrow (5 Th), DIA de novo sequencing significantly outperforms DDA, identifying almost twice as many peptides. This suggests that in older-generation mass spectrometers, smaller isolation windows in DIA mode allow for more precise detection, enhancing the depth of peptide identification. However, as the isolation window increases to 10 Th and 20 Th, the performance of DIA begins to diminish in both de novo sequencing and database search, losing its advantage over DDA. At a 20 Th isolation window, while DIA database search still detects more peptides than DDA, de novo sequencing lags behind. The additional peptides identified in database search mode cannot be accurately recovered by de novo sequencing. This decline illustrates that larger isolation windows introduce higher-level of coelution, leading to unwanted noise and overlapping ion signals, reducing the accuracy and efficiency of peptide detection. Therefore, increasing the isolation window in older-generation mass spectrometers appears to be a suboptimal approach for DIA experiments.
In contrast, the performance of the Orbitrap Astral mass spectrometer, using the PXD046453 [32] HEK293T dataset, presents a more consistent scenario (shown in Figure 4) due to the enabled narrow window mode. This dataset provides both DDA data and DIA data acquired at different cycle times (2.5 ms and 3.5 ms), both using a narrow 2 Th isolation window. The Astral data demonstrates that DIA consistently outperforms DDA, regardless of the cycle time. The DDA de novo sequencing identifies approximately 17,000 peptides, whereas the DIA de novo sequencing detects nearly 32,000 peptides, showing a considerable increase over DDA. The additional  15,000 peptides identified by DIA de novo are likely missed by DDA due to biased sampling. This highlights the key advantage of DIA over DDA—its comprehensive and unbiased sampling of peptides, which is further enhanced by the speed and sensitivity of the Orbitrap Astral. In this case, the smaller isolation window and improved performance of Astral’s DIA mode enable the identification of a broader range of peptides, overcoming the limitations observed with DDA.
Overall, the results of both datasets underscore the advantages of using DIA de novo sequencing with narrow isolation windows, particularly when coupled with the advanced capabilities of next-generation instruments like the Orbitrap Astral. This mode not only boosts peptide detection rates but also minimizes the sampling bias inherent in DDA, providing a more comprehensive view of the proteome.
2.4 Comparison with Cascadia [21]
To evaluate the performance of our approach in a practical setting, we conduct a comparative analysis against Cascadia, a recently introduced state-of-the-art model. For this purpose, we selected a representative raw file from the PXD046386 Yeast Knockout (KO) dataset. This dataset provides a relevant benchmark for assessing model effectiveness in proteomics data analysis. Both our method and Cascadia were applied to this identical data source under equivalent preprocessing and parameter settings to ensure a fair comparison. The resulting performance differences are summarized and visualized in Figure 5.
Figure 5: Venn diagram for Cascadia comparison.
Our model demonstrates a 21% increase in peptide identifications compared to Cascadia. Notably, the overlap in identified peptides between the two models is relatively limited, indicating that each model captures distinct subsets of the data. This suggests that our approach exhibits different identification characteristics, potentially offering complementary insights to those provided by Cascadia.
2.5 Ablation Study
In this ablation study, illustrated in Figure 6, we evaluate the impact of different configurations on our model using the Yeast KO dataset. The results highlight that coelution-aware pretraining plays a helpful role in model performance, as its removal reduces peptide recall by 8.5% compared to DIANovo. More notably, the removal of Rotary Positional Encoding (RoPE) leads to a severe performance drop, with peptide recall plunging to 67.2% of the original. This decline suggests that absolute positional embeddings alone are insufficient for encoding graph node masses effectively. Additionally, excluding extended peak features or limiting the input to only three neighboring spectra results in moderate performance losses, demonstrating the importance of both extended features and broader spectral context. Finally, binning spectra into 0.5 Th intervals causes a significant drop in performance due to reduced spectral resolution, which increases ambiguity in peptide identification.
((a)) Relative amino acid recall
((b)) Relative peptide recall
((c)) Relative amino acid precision
Figure 6: Relative amino acid recall (a), peptide recall (b), and amino acid precision (c) of our method vs different configurations, compared to our method, on the Yeast KO dataset.
2.6 Theoretical Analysis of De Novo Peptide Sequencing Performance
We present a theoretical framework to explain the performance variations observed in de novo peptide sequencing across different mass spectrometry acquisition methods: Data-Dependent Acquisition (DDA), older-generation Data-Independent Acquisition (DIA), and Astral DIA. Our analysis focuses on how the balance between signal and noise in the generated spectra affects the efficacy of peptide matching algorithms like XCorr in both database searches and de novo sequencing.
  • Signal and Noise Characteristics in Different Acquisition Methods
    We define signal peaks as the fragment ions originating from the target peptide. All other peaks are considered noise, which typically includes fragment ions from coeluting peptides, immonium ions, and instrumental noise. In DDA spectra, we typically observe approximately 50 signal peaks corresponding to fragment ions of the peptide of interest, along with about 500 noise peaks. Transitioning from DDA to older-generation DIA results in a slight increase in signal peaks to around 60 but introduces a substantial increase in noise peaks to approximately 1,750, due to spectra being highly-multiplexed. We argue that this significant escalation in noise outweighs the marginal gain in signal peaks, leading to diminished de novo sequencing performance in older-generation DIA compared to DDA.
    When moving from older-generation DIA to Astral DIA, one might expect the scenario to revert to that of DDA due to the employment of narrower isolation windows. Contrary to this expectation, Astral DIA produces a considerable increase in signal peaks to approximately 150 and an even larger surge in noise peaks to around 9,000, while demonstrating high level of coelution in spite of the narrow isolation window. Additionally, Astral DIA data exhibits lower noise peak intensity relative to signal peaks compared to older-generation equipment. Despite the higher noise levels, Astral DIA demonstrates significantly better de novo sequencing performance than older-generation DIA.
  • Theoretical Model Explaining the Observed Performance
    To elucidate this phenomenon, we propose a theoretical model based on the difficulty of matching peptides against noise using the widely adopted peptide matching algorithm, XCorr. Our central argument is that a higher XCorr value for a de novo peptide indicates a more straightforward and confident match of the peptide sequence in both database searches and de novo sequencing algorithms.
    In database searches, the XCorr score [33] quantifies the similarity between the experimental spectrum and theoretical spectra generated from candidate peptides in the database. A higher XCorr score reflects a better match, leading to increased confidence in peptide identification. This is because the probability of a high-scoring match occurring by chance decreases exponentially with increasing score, enhancing the specificity of the identification.
    For de novo sequencing, the advantage of a higher XCorr score stems from the intrinsic similarities between database search algorithms and de novo methods. De novo algorithms are heuristic approaches designed to reconstruct peptide sequences directly from spectra without relying on a predefined database. Although de novo methods may employ advanced scoring functions tailored for sequence reconstruction, the XCorr score serves as a valuable lower-bound approximation of their performance. A higher XCorr score implies that the spectrum contains clear and informative fragment ion peaks that can be effectively utilized by de novo algorithms to deduce the peptide sequence.
  • Simulation of Experimental Scenarios
    To validate our theory that signal-noise profile plays a crucial rule in de novo performance, we conducted simulations under the three experimental conditions: older-generation DDA, older-generation DIA, and Astral DIA. We generated synthetic spectra that reflect the characteristic signal and noise levels associated with each of the three methods. Additionally, we created synthetic spectra that replicate the signal and noise profiles of each test dataset, accounting for peptide length and noise intensity, thereby simulating real experimental conditions. Utilizing the approach proposed by Noble et al. (2012)[34], we computed the p-value for a single peptide-spectrum match based on dynamic programming and XCorr scoring. This method allows for the estimation of the statistical significance of the observed XCorr scores, providing insight into the confidence of peptide identifications under varying signal-to-noise conditions.
  • Relationship Between p-value and Peptide Recall
    In theory, the p-value should exhibit an inverse correlation with peptide recall, as discussed in detail in Section 4.6. This theoretical relationship is supported by our simulation results using real test datasets, as shown in Figure 7. The figure clearly demonstrates an inverse relationship between peptide recall and p-value, indicating that lower p-values are associated with higher recall rates. Furthermore, the results reveal a clear separation between older-generation DIA and Astral DIA data. Older-generation DIA data predominantly occupies the upper-left quadrant, characterized by lower peptide recall and higher p-values. In contrast, Astral DIA data is concentrated in the lower-right quadrant, reflecting its higher peptide recall and lower p-values. This distinction highlights the superior performance of Astral DIA, which can be attributed to its improved signal-to-noise profile and lower noise intensity, resulting in enhanced data quality and spectral characteristics.
    This finding highlights the utility of the simulated p-value as a reliable indicator of de novo sequencing performance. By effectively capturing the underlying relationship between statistical significance and identification accuracy, the p-value provides valuable insights into the quality of peptide identifications. These results further validate the robustness of our simulations and their alignment with real experimental conditions, reinforcing the practical relevance of this theoretical framework.
    Figure 7: Simulated p-values vs peptide recall for test datasets, with a Pearson correlation coefficient of -0.68.
Our simulation results are presented as a heatmap in Figure 8, illustrating the impact of the signal-to-noise profile on de novo sequencing performance. The three experimental scenarios are labeled within the figure for reference. The simulations reveal that the computed p-values align with our theoretical expectations. Specifically, both older-generation DDA and Astral DIA exhibit similar p-values, which are significantly lower than those observed for older-generation DIA. This finding suggests that peptide-spectrum matching is intrinsically more favorable in the contexts of Astral DIA and DDA compared to older-generation DIA. The improved performance of Astral DIA compared to older-generation methods demonstrates that, despite its higher noise levels, it also produces significantly more signal peaks. In this scenario, the increased number of signal peaks effectively compensates for the additional noise. This suggests that the absolute signal peak count plays a more critical role than the signal-to-noise ratio, as Astral DIA yields a lower signal-to-noise ratio than older-generation methods yet still achieves superior results.
Figure 8: Simulated p-values for different signal and noise values. In the figure, older-generation DDA points to 50 signal peaks and 500 noise peaks with p-value of 0.00591, older-generation DIA points to 60 signal peaks and 1750 noise peaks with p-value 0.01144, while Astral-DIA points to 90 signal peaks and 9000 noise peaks with p-value 0.00516.
3 Discussion
This study introduces a robust and highly accurate method for de novo peptide sequencing within the DIA setting, offering substantial improvements over existing approaches. By incorporating dilated convolutional neural networks, FlashTtention-2, RoPE, as well as coelution-aware pretraining, our model is adept at handling the complex, highly-multiplexed and noisy spectral data inherent to DIA experiments. The use of these advanced techniques allows our approach to effectively capture intricate spectral patterns, enabling reliable peptide sequence predictions even in challenging conditions.
Our comprehensive experiments demonstrate the superior performance of the proposed method, particularly when applied to data from the Orbitrap Astral mass spectrometer. The Astral’s enhanced fragment ion coverage and consistent fragmentation patterns provide a solid foundation for our de novo sequencing approach, leading to a marked increase in peptide identification accuracy, even in case of high coelution level. In contrast, older-generation mass spectrometers, such as those used in older DDA and DIA experiments, reveal the limitations of existing de novo sequencing methodologies, especially when larger isolation windows are used. Our DDA vs. DIA comparison study highlights these challenges, showing that while DIA outperforms DDA with narrower isolation windows, it begins to lose this advantage as the window size increases. This issue is particularly evident with older-generation instruments, where DIA performance in both database search and de novo sequencing diminishes at larger window sizes. However, the Orbitrap Astral consistently demonstrates DIA’s superiority over DDA, facilitated by the narrow isolation windows enabled by this instrument, underscoring the pivotal role that modern mass spectrometers play in advancing peptide sequencing capabilities. In addition, our ablation study verifies the validity and neccessity of our design.
During our theoretical analysis of de novo performance, we show that the balance between signal enhancement and noise accumulation is critical. In the case of Astral DIA, the substantial increase in informative signal peaks provides a richer dataset for de novo algorithms to process, improving the accuracy and confidence of peptide sequence reconstruction. This outcome underscores the importance of optimizing acquisition parameters to maximize signal quality while managing noise levels.
Our findings suggest that acquisition methods like Astral DIA, which can substantially increase signal peaks while effectively managing noise, are more conducive to de novo peptide sequencing. This has important implications for the development of mass spectrometry techniques and the selection of acquisition parameters in proteomics research. Conversely, older-generation DIA does not offer the same advantage because the modest gain in signal peaks is insufficient to counterbalance the significant escalation in noise peaks. This results in lower confidence in peptide identifications, as reflected by higher p-values.
In conclusion, our study represents an advancement in de novo peptide sequencing within the DIA framework. By integrating advanced computational techniques to tackle the coeluted nature of DIA spectra, and leveraging the advanced capabilities of next-generation mass spectrometers like the Orbitrap Astral, we offer a robust and accurate solution for analyzing complex proteomic data. The findings of this study not only demonstrate the limitations of older-generation instruments in de novo peptide sequencing but also highlight the potential of DIA to provide more comprehensive and precise peptide identification. This work paves the way for future innovations in de novo sequencing, offering new opportunities for discovering novel peptides and advancing our understanding of proteomic diversity.
4 Methods
Our model architecture is an encoder-decoder Transformer operating on the spectrum graph, including spectrum graph construction, dilated CNN (convolutional neural network) layers to process the chromatograms, RoPE (rotary positinal embedding)[40] integration to encode mass differences in the graph, coelution-aware pretraining to provide better understanding of the multiplexed spectra, and the two-stage decoding, where we predict the optimal path through the spectrum graph in stage 1, and refine the path to generate the final peptide sequence with mass tag filling in stage 2, in order to alleviate the sequence memorization problem commonly associated with de novo sequencing algorithms, where the model relies too much on previous seen peptide sequences during training. The model architecture is shown in Figure 9.
Figure 9: The model structure of Our entire workflow. On the top is the optimal path task, generating a series of node indices, which are transformed into the optimal path. The mass values in the optimal path are then translated to the corresponding amino acids when a single match is found. On the button is the sequence generation task. It takes the generated optimal path as input and outputs the amino acid sequence to replace mass tags. In the figure, FFNGLU refers to Feed-Forward Network with Gated Linear Units, commonly used in Transformers and deep learning architectures to enhance expressiveness and efficiency. CNN refers to Convolutional Neural Networks, and RoPE refers to Rotary Positional Embeddings.
4.1 Feature Extraction
In a tandem mass spectrometry experiment, peptide precursors are first analyzed in their intact form, generating first-stage scans (MS1 scans). The peptides are then fragmented, and the resulting fragments are analyzed again, producing second-stage scans (MS2 scans).
The feature construction for the encoder is designed to capture the chromatogram characteristics in DIA, in order to better distinguish signal peaks from coeluting fragment ions and noise. For each PSM (peptide-spectrum match), we collect five neighboring MS2 spectra along the retention time (RT) dimension, centered around the RT peak. This selection is crucial as it differentiates DIA data from DDA. The neighboring spectra provide a chromatogram for each observed peak in the spectrum, and the relationship among these chromatograms improves the robustness of peptide sequence predictions. [29]
For MS1 spectra, alignment with corresponding MS2 spectra is needed. To achieve this, MS1 intensity values are linearly interpolated based on their RT values to match the MS2 spectra exactly, ensuring that the features derived from MS1 and MS2 spectra are directly comparable.
During preprocessing, each spectrum is binned into fixed-size 0.01 Th intervals, enabling the construction of chromatograms for each m/z value. Once the chromatograms are generated through binning, the spectra are reconstructed as a list of (m/z, chromatogram) pairs for neural network processing.
In addition to the primary spectral data, our methodology involves the calculation of eight supplementary features to enrich the feature set used for encoding, adapted from GraphNovo[13] and Novor[41]. These features are derived from each spectrum within the ±$\pm 50$ Th neighboring window and are designed to capture various statistical and intensity-based characteristics of the spectral peaks.
Post feature extraction, we compile the data into two structured tensors for each m/z value in combined spectrum:
  • A $[5,1]$ tensor representing the chromatogram data across the five neighboring MS2 spectra.
  • A $[5,8]$ tensor that encapsulates the eight computed features across the same neighboring spectra.
Finally, the spectrum graph is constructed by transforming every peak in the original spectrum from m/z to N-terminal residual masses. Each peak in the original spectrum is converted into 6 peaks in spectrum graph, taking into account 6 types of ions, including $a^{+}$, $a^{2+}$, $b^{+}$, $b^{2+}$, $y^{+}$, and $y^{2+}$. We operate on the N-terminal residual masses, denoted as the graph node mass, following Equation 1, during which C-terminal ions are converted to their N-terminal residual masses by subtracting their C-terminal residual mass from the precursor mass. Note that although we did not specifically include other common ion types like $b-H_{2}O$ or $y-{NH_{3}}$, their peaks are converted into graph node as well. They do not belong to the target sequence, but their graph node can still contribute to our de novo objective by self attention. We provide an example of spectrum graph in Figure 10.
Figure 10: An example of a spectrum graph is shown, where the bottom value on each edge represents the mass difference between nodes, encoded by RoPE, and the top value indicates the corresponding amino acid sequence. Only a subset of nodes and edges is plotted for clarity, whereas, in a complete spectrum graph, all possible forward connections would be present.
$m_{\text{nterm}}=\begin{cases}\left(m/z-m_{\text{proton}}\right)\times c-\sigma,&\text{if ion}\in\{a,b,c\text{-ions}\}\\ m_{\text{precursor}}-\left(m/z-m_{\text{proton}}\right)\times c+\sigma,&\text{if ion}\in\{x,y,z\text{-ions}\}\end{cases}$ (1)
where offset $\sigma$ depends on the ion type, computed as Supporting Table 1; $c$ is the charge of ion; $m_{proton}$ is the mass of proton and $m_{precursor}$ is the precursor mass [13].
4.2 Time-Series Encoder and Spectrum Encoder
In our approach, we employ a dilated convolutional neural network (CNN) to effectively encode the time-series data derived from the chromatogram and the additional spectral features.
Dilated CNNs [42] are a variant of traditional convolutional networks. The primary distinction of dilated convolutions lies in their ability to expand the receptive field without losing resolution or coverage, achieved by inserting gaps (known as dilation) between each element in the convolution kernel. This allows the network to encode time-dependent information in chromatograms while keeping the number of parameters and memory/computational complexity relatively low.
For the chromatogram and its associated spectral features, dilated CNNs are ideally suited. Our model processes inputs structured as $[5,1]$ or $[5,8]$ tensors, where ‘$5$’ represents the number of neighboring spectra considered, ‘$1$’ denotes the chromatogram, and ‘$8$’ reflects the additional computed features. By applying dilated convolutions, the model can integrate both local and broader contextual information across these inputs. The outputs of the dilated CNN are encoded into tensors of size $[hidden\_size]$, where ‘$hidden\_size$’ indicates the dimension of the feature space. This encoding captures information about the temporal dynamics of a specific m/z value.
The time-series embeddings are input directly into the Transformer spectrum encoder, where the self-attention mechanism learns to identify similarities in the time-series data. Ideally, ions from the same peptide will exhibit similar chromatograms, whereas ions from coeluting peptides will display different patterns. By explicitly modeling these similarities, the model’s ability to differentiate between ions from coeluting peptides is improved.
In addition, we adopt the FlashAttention 2 [26] implementation of the attention function, allowing us to process very large DIA spectra, with high computation and memory efficiency.
4.3 RoPE Integration
GraphNovo employs the Graphformer[43] graph neural network to encode the spectrum graph, capturing comprehensive information about the spectrum. However, applying this approach to DIA data is impractical due to the large size of DIA spectra, which results from high levels of coelution. To address this challenge, we replace the memory-intensive node and path encoders used in Graphformer with Rotary Position Embedding (RoPE). This step is also neccessary for the adoption of Flash Attention 2, since it does not allow us to operate directly on the attention matrix, which is required by Graphformer.
RoPE is a position encoding technique that introduces rotational invariance by representing positional information in a circular format[40]. In our approach, RoPE encodes the mass differences between graph nodes, allowing the model to learn meaningful mass relationships automatically. This effectively transforms the spectrum graph into a fully-connected graph where edges represent mass differences, enabling the model to identify the most relevant mass differences for the task without the computational overhead of traditional node and path encoders. The adoption of RoPE is advantageous, as it captures the critical information conveyed by the mass difference between pairs of graph nodes, effectively representing the cumulative mass of the amino acids that lie between them.
4.4 Two Stage Decoder
We adopt a two-stage decoding process, similar to GraphNovo[13] to alleviate the sequence memorization issue during training, where the decoder simply remembers seen training sequences, and fail to generalize on unseen sequences. This approach includes:
  • Stage One: Optimal Path Prediction
    In the first stage, the model predicts the optimal path through the spectrum graph. This involves identifying the most probable sequence of nodes (representing graph node mass values) that form a potential peptide sequence. The graph is constructed such that each node represents a possible peptide fragment, and edges denote feasible transitions based on mass differences.
  • Stage Two: Sequence Filling
    In the second stage, the optimal path is refined to generate the final peptide sequence. This stage involves filling in the mass tags along the predicted path with corresponding amino acids, ensuring the sequence adheres to known peptide fragmentation patterns.
Readers can refer to Figure 1a of Mao et al. (2023)[13] for a detailed visualization of the two-stage decoding process.
4.5 Coelution-aware Pretraining
A key difference between DDA and DIA data is that, in DDA, since precurosr selection is performed, there is typically a small number of peptides in the MS2 spectrum. Meanwhile in DIA, since we uniformly select a precursor range to perform fragmentation, there is a significant amount of coelution, i.e., one spectrum consisting of fragment ions of multiple coeluting peptides. In our study, coeluing peptides are identified by searching through the database search result, and include those whose RT (retention time) overlaps with the target de novo peptide, and their precursor m/z falls in the same isolation window as the target peptide.
In a normal de novo sequencing algorithm, we typically only consider the fragment ions of the de novo target peptide. For instance, in the optimal path task of our program, we predict only the graph nodes (n-terminal masses of peaks) belonging to the de novo target peptide, and the next step we replace these graph nodes with amino acids, resulting in predicted peptide sequence.
However, if we incorporate the information about the fragment ions of other coeluting peptides, we might be able to provide the model broader information about the spectrum, improving de novo performance. More specifically, in traditional de novo algorithms, the fragment ions of other coeluting peptides are treated as noises, but we can give them labels for the model to learn. With such information, the model can make less mistakes distinguishing noise peaks from signal peaks (of the target de novo peptide), since the model knows some noise peak is probably a fragment ion of other coeluting peptide, thus less likely predict it as a signal peak.
To achieve this, we introduce coelution-aware pretraining to our algorithm. We adopt the same Transformer encoder in Section 4.1, without the spectrum graph conversion, i.e. we keep all the original m/z values as the input to the Transformer. After layers of self-attention, we obtain the embedding for each m/z value in the spectrum. These embeddings are trained under the ion type loss, a cross entropy loss, representing the type of fragment ions. 12 types of ions are considered (see Supporting Table 2), plus noise. These labels corresponds to the fragment ion types of all coeluting peptides, not only the target peptide.
The workflow for the pretraining model is illustrated in Figure 11. After pretraining, the trained embeddings are fed to downstream optimal path and sequence generation models as features for the Transformer encoder.
Figure 11: Pretrain model architecture.
4.6 Theoretical Analysis
We implemented simulations to generate synthetic mass spectra reflecting the characteristic signal and noise levels associated with DDA, older-generation DIA, and Astral DIA methods.
  • Addition of Signal and Noise Peaks
    Signal peaks were uniformly sampled from the theoretical peaks of the target peptide to represent peptide evidence. To simulate experimental conditions, random noise peaks were added to each spectrum, with the quantity adjusted to reflect typical noise levels for each acquisition method, introduced by different level of coelution. Specifically, for DDA, we included 50 signal peaks and 500 noise peaks; for older-generation DIA, 60 signal peaks and 1,750 noise peaks were added; and for Astral DIA, 150 signal peaks and 9,000 noise peaks were included. Noise peaks were assigned random m/z values with uniform intensities to mimic realistic spectral conditions.
    For real world test dataset, we substitute the experimental parameters with real data statistics, and take into consideration the median peptide length, as well as median signal and noise intensities.
  • Calculation of XCorr Scores
    We calculated the XCorr scores for the simulated spectra by cross-correlating the experimental spectra with the theoretical spectra of the target peptides. This involved binning the spectra, normalizing the intensities, and computing the dot product between the experimental and theoretical spectra after shifting the theoretical spectrum over a range of lags to find the maximum correlation.
    Computation of p-values: We employed the dynamic programming approach described by Noble, et al. [34] to compute the p-values associated with the XCorr scores. This method estimates the distribution of XCorr scores under the null hypothesis and determines the statistical significance of the observed XCorr scores.
  • Adjustment for Multiple Hypotheses
    Theoretically, the p-value is inversely related to peptide recall. This inverse relationship arises because, when computing the Sidak-corrected p-value [44], we must adjust for a significantly larger number of peptide hypotheses in de novo sequencing than in database searches. In database searches, only the peptides existing in the database are considered, limiting the number of comparisons. In contrast, de novo sequencing must account for all possible peptide sequences, dramatically increasing the number of hypotheses and necessitating a stricter correction for multiple testing. Consequently, the p-value threshold becomes more stringent in de novo sequencing, potentially reducing peptide recall. The relationship between peptide recall and p-value is given by the following equation:
    $Peptide\ Recall=\mathbb{E}{\left[\frac{\#Denovo\ Success}{\#Database\ Success}\right]}\approx\frac{(1-p)^{\#\ De\ Novo\ Peptides}}{(1-p)^{\#\ Database\ Peptides}}$ (2)
    where #$\#\ De\ Novo\ Success$ and #$\#\ Database\ Success$ are the number of successfully identified peptides by de novo mode/ database search mode, $p$ is the p value, and #$\#\ De\ Novo\ Peptides$ and #$\#\ Database\ Peptides$ refer to the number of peptides with the same precursor mass to compare for de novo or database search, with #$\#\ De\ Novo\ Peptides\gg\#\ Database\ Peptides$
5 Conclusion
In this paper, we present a novel and highly effective method for de novo peptide sequencing in the DIA setting, offering significant improvements over traditional approaches. A key focus of our study is the comparison between DDA and DIA, where we demonstrate that while DIA outperforms DDA with narrow isolation windows, its advantage diminishes with wider windows in older-generation instruments. However, with the Orbitrap Astral, DIA consistently surpasses DDA, highlighting the importance of next-generation mass spectrometers in improving peptide identification. To explain this phenomenon, we propose a simulation based theory.
Our method, specifically designed to tackle the highly mutliplexed DIA data, along with leveraging advanced instrumentation, provides robust performance in challenging proteomic environments. These findings emphasize the value of using DIA with modern technologies for comprehensive peptide detection, paving the way for more reliable and effective de novo sequencing methods in proteomics.
Appendix A Additional Features for Spectrum
Below is a list of the 8 additional features about the spectrum in our model.
  1. 1.
    Normalized Mass Over Charge: We compute the mass-to-charge ratio, normalized by a predefined upper limit of $3500Th$, to standardize this value across different spectra. Computed as $e^{\frac{m/z}{UPPER\_LIMIT}}$.
  2. 2.
    Relative Intensity: Each peak’s intensity is expressed as a fraction of the intensity of the most abundant peak in the spectrum, facilitating comparisons across variable signal strengths. Computed as $\frac{I}{I_{max}}$, where $I_{max}$ is the intensity of most abundant peak in the spectrum.
  3. 3.
    Rank: Each peak is assigned a rank based on its abundance relative to other peaks, sorted from the most to the least abundant. Computed as $\frac{R}{N}$, where $R$ is the target peak’s rank in terms of intensity.
  4. 4.
    Half Rank: This metric assigns a rank to a peak after the intensities of all peaks are halved, highlighting the relative stability of peak abundances. Computed as $\frac{R_{half}}{N}$, where $R_{half}$ is the target peak’s rank with half its intensity.
  5. 5.
    Local Significance: Using the hyperbolic tangent function, we scale the intensity of each peak relative to the minimum intensity within the neighboring window, emphasizing peaks with significant local variance. Computed as $tanh({\frac{I}{2(I_{min}-1)}})$, where $I_{min}$ is the lowest intensity within 50 Th window.
  6. 6.
    Local Rank: Similar to the global rank, but restricted to peaks within the ±$\pm 50$ Th m/z range, providing a localized perspective on peak significance. Computed as $\frac{R_{local}}{N_{local}}$, similar to rank.
  7. 7.
    Local Half Rank: This is computed like the half rank but limited to the local window, offering insights into the comparative dynamics of local peaks. Computed as $\frac{R_{half\_local}}{N_{local}}$.
  8. 8.
    Local Relative Intensity: We measure each peak’s intensity relative to the most intense peak within the neighboring window, allowing for an assessment of local peak dominance. Computed as $\frac{I}{I_{local\_max}}$.
Appendix B Ion Offsets
Supporting Table 1 includes the ion offsets for converting to n-terminal mass in our model.
Ion Type Neutral Offset
a [N] - CHO
b [H] - H
y [C] + H
List of supportingtables 1 Ion offsets for ion types we used. [N] is the molecular mass of the neutral N-terminal group; [C] is the molecular mass of the neutral C-terminal group. C,H,O are the mass of the carbon atom, hydrogen atom and oxygen atom individually. [13]
Appendix C Type of Labeled Ions in Pretraining
Supporting Table 2 includes the ion types considered in our pretrain model.
Ion Type
1a
1b
2a
2b
1a-NH3
1a-H2O
1b-NH3
1b-H2O
1y
2y
1y-NH3
1y-H2O
List of supportingtables 2 Types of Ions Labeled during coelution-aware pretraining. For precursor charge $\leq$ 2, ions with charge 1 are considered. For precursor charge > 2, ions with charge 1 or 2 are considered.
Appendix D Precursor Feature Detection
For the detection of precursor features from liquid chromatography-mass spectrometry (LC-MS) maps, we utilized the same standard set of precursor information as employed by DeepNovo-DIA [19]. In our experiments, we implemented the detection results from DIA-NN [29] during the main experiment, and DIAUmpire [45] during the comparison with Cascadia; however, these can be substituted with outputs from other existing peak detection algorithms, such as those referenced in Zhang et al. (2012) [38], Taynova et al. (2016) [46] or Tsou et al. (2015) [45]. The outcome of this detection step is a list of precursor features, each comprising the following essential information: feature ID, precursor mass-to-charge ratio (m/z), charge state, retention-time center, and scans across the retention-time range.
Furthermore, given the m/z and retention-time range of a feature, we collected all tandem mass spectrometry (MS/MS) spectra that fell within the feature’s retention-time range and whose DIA m/z windows encompassed the feature’s m/z value. More specifically, our precursor information includes the following data:
  • Feature ID: an unique identifier given to each precursor.
  • Precursor m/z: the mass-to-charge ratio of the precursor ion.
  • Precursor Charge: the charge state of the precursor ion.
  • RT_Mean: the mean of the retention-time range.
  • Sequence: this column remains empty during de novo sequencing; in training mode, it contains the peptide sequences identified by the in-house database search for training purposes.
  • Scans: a list of all MS/MS spectra associated with the feature as described above.
Appendix E Characteristics of Test Datasets
Supporting Table 3 provides the key parameters for each test dataset, these parameters are utilized in our theoretical analysis of peptide recall.
Dataset Coeluting Number # Signal Peaks # Noise Peaks Median Peptide Length Median Noise Intensity Isolation Window Mass Spectrometer
Hela 5 Th 18.50 72 1,965 11 0.44 5 Q Exactive
Hela 10 Th 28.03 80 1,965 12 0.52 10 Q Exactive
Hela 20 Th 36.20 91 2,438 13 0.56 20 Q Exactive
PXD026600 Narrow 5.73 59 967 13 0.64 8 Fusion
PXD026600 Wide 7.24 65 1,485 14 0.69 15 Fusion
PXD026600 Overlap 5.79 53 1,130 13 0.65 8 Fusion
PXD026600 Variable 5.95 60 1,275 13 0.67 8-15 Fusion
PXD019777 12.34 91 3,057 13 0.73 24.25 Q Exactive
Yeast KO 19.49 152 9,404 12 0.34 3 Astral
PXD046453 2p5ms 10.15 152 12,284 12 0.21 2 Astral
PXD046453 3p5ms 11.99 155 14,069 12 0.18 2 Astral
PXD046283 100ng 4.66 82 2,790 12 0.54 2 Astral
PXD046283 300ng 5.01 105 4,007 12 0.49 2 Astral
PXD046283 1000ng 5.41 131 6,114 12 0.40 2 Astral
PXD046283 2000ng 6.04 137 6,960 12 0.37 2 Astral
PXD046471 30min 6.97 136 6,526 12 0.34 2 Astral
PXD046471 7min 14.04 130 6,359 12 0.51 4 Astral
List of supportingtables 3 Experimental Parameters for test datasets. Coeluting number refers to how many coeluting peptides one peptide has on average, number of signal or noise peaks refers to the median number of peaks of the neighboring five spectrums which are fragment ions or the target peptide or not, median noise intensity refers to the ratio between median intensity for noise peaks and signal peaks, and isolation window has unit Th.
Appendix F Links to Datasets
Supporting Table 4 provides links to our training and testing datasets. Every dataset is searched with DIA-NN 1.8.1 [29] with carboxymethylation of cysteine as fixed modification, and oxidation of methionine as variable. PSMs under 1% FDR is kept for training or testing. The raw files are converted to mzML with msConvert 3.0 with vendor peak picking. Then the PSMs and mzMLs are converted to our customized format with our processing codes, by extracting the related spectrums of each PSM. For de novo only analysis, feature detection is done on the mzML files with DIA-Umpire 2.2.8 [45].
List of supportingtables 4 Links to each dataset
Appendix G Model Implementation Details
Our model is trained on a Lion optimizer [49] with learning rate $2\times 10^{-6}$, over 3 epochs, with early stop based on validation error. The model includes 4 layers both on encoder and decoder side, with 1,024 hidden size and 8 attention heads.
For older-generation data, training and validation sets are Pain, PXD019777, and PXD003179, with 680,947/254,467/1,206,052 PSMs respectively. For Astral data, they are PXD046386 or PXD046453 + PXD046453, with 1,896,662, 1,086,577, and 2,136,215 PSMs respectively, these PSMs are chosen from the DIA-NN search result with 1% FDR.
To prevent information leakage and ensure a rigorous evaluation of model generalization, we designed our training/validation/test splits to fully separate related peptide signals across sets, including coeluting peptides. Our model architecture leverages not only the target peptide but also contextual information from coeluting peptides within a spectrum. Therefore, to eliminate any direct or indirect sequence overlap, we adopted a raw file–based splitting strategy: for each training run, we designated a specific RT range in one raw file exclusively for validation and used the remaining for training. This guarantees that no peptides, or their coeluting partners, appear in both training and validation sets.
Each training scheme was associated with a held-out validation set comprising 20,000 peptide-spectrum matches (PSMs). These validation examples were used strictly for hyperparameter tuning and early stopping. No validation data was used in model selection or test-time inference.
For the test set, we ensured strict non-overlap in both peptide sequences and experimental conditions. All test sequences were drawn from datasets entirely distinct from those used in training. Furthermore, in our cross-species evaluation using Orbitrap Astral data, we imposed an even stricter criterion: training and testing data come from different species. Specifically, when training on the human datasets PXD046453 + PXD046444, we tested only on the yeast dataset PXD046386. Conversely, when training on the yeast dataset PXD046386, we used entirely separate human datasets for testing.
Appendix H Visualization of Learned Embeddings
In this section, we present a visualization of the learned peak embeddings to better understand the features captured by our pretrained model.
We generate a t-SNE [50] plot from the learned peak embeddings, and observe that the model implicitly captures the relationship between coeluting precursors and their corresponding fragment ions, even when trained solely with the ion type loss. The t-SNE plots (shown in Supporting Figure 1, 2, and 3) further reveal that peak embeddings are organized not only by fragment ion type (ion type label) but also by their source peptide (ion source label). For instance, in Supporting Figure 1, a dark blue cluster on the right side of the graph represents peaks originating from peptide 1 in the ion source plot. Within this cluster, the peaks are further subdivided into distinct colors (pink and green) in the ion type plot, corresponding to different fragment ion types. We provide the t-SNE plots from three different peptides to illustrate this effect.
These results suggest that the model inherently learns to associate each fragment ion with its coeluting peptide, highlighting its ability to extract meaningful structural relationships from the data.
((a)) Ion Type Label
((b)) Ion Source Label
List of supportingfigures 1 tSNE plot for ion type label (a) and ion source label (b) of peptide DHGEGGIIVGSALENK2. The color for ion type label refers to the fragment ion type of each peak, while the color for ion source label refers to which coeluting peptide a peak comes from. Noise peaks (which do not belong to any coeluting peptide) are excluded.
((a)) Ion Type Label
((b)) Ion Source Label
List of supportingfigures 2 tSNE plot for ion type label (a) and ion source label (b) of peptide GYWGTNLGQPHSLATK2.
((a)) Ion Type Label
((b)) Ion Source Label
List of supportingfigures 3 tSNE plot for ion type label (a) and ion source label (b) of peptide EYLPEMAASYSHPK2.
Appendix I Sensitivity to Coelution Number
In this section, we investigate how our algorithm responds to varying levels of coelution. The coelution number is defined as the number of peptides that coelute with a target peptide, specifically when their precursor m/z values fall within the same precursor isolation window and their retention times overlap.
To visualize the relationship between coelution number and de novo sequencing performance, we generated a plot (shown in Supporting Figure 4) that demonstrates our algorithm’s performance across different levels of coelution. The plot illustrates that our algorithm maintains consistent performance irrespective of the coelution number, indicating its robustness in handling complex spectral data.
List of supportingfigures 4 Performance of our method vs baselines, on the Yeast KO Dataset, on peptides with varying coelution number.
The ability of our method to maintain performance in the presence of numerous coeluting peptides underscores its effectiveness in dealing with the inherent complexity of DIA data. This resilience is critical for practical applications in proteomics, where accurate peptide detection amidst complex mixtures is essential.